40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 18     Entry time: Fri Oct 26 16:19:29 2007
Author: Tobin Fricke 
Type: Routine 
Category: IOO 
Subject: MC resonances 
We would like to measure the absorption of the mode cleaner optics. The plan is to repeat <a href="http://ilog.ligo-wa.caltech.edu:7285/mLIGO/Cleaning_the_Mode_Cleaner">Valera's experiment</a> in which we track the MC's thermal resonances to infer their power absorption. Last night Rana and I hooked up a lock-in amplifier to heterodyne the MC servo signal by 28 kHz and piped the output into an ADC using the MC_AO channel. We did not find any resonances.

Valera recommends we drive the POS of the three MC optics with bandlimited noise to excite the resonances.
ELOG V3.1.3-